OpenStudy (anonymous):

A street that is 156m long is covered in snow. City workers are using a snowplow to clear the street. A tire on the snowplow has to turn 39 times in traveling the length of the street. What is the diameter of the tire? Use the value 3.14 for Pi . Round your answer to the nearest tenth. Do not round any intermediate steps.

OpenStudy (saifoo.khan):

156/39

OpenStudy (bahrom7893):

wrong

OpenStudy (saifoo.khan):

That will give you it's circumfrence.

OpenStudy (bahrom7893):

correct

OpenStudy (anonymous):

around 13?

OpenStudy (anonymous):

I lied...

OpenStudy (anonymous):

i got 4

OpenStudy (saifoo.khan):

Then, use the formula, Cirmufrence = pi * d Where d is the diameter.

OpenStudy (saifoo.khan):

4/3.142 = d

OpenStudy (anonymous):

nvm I didn't... 12.4

OpenStudy (anonymous):

^ close enough to 13

OpenStudy (anonymous):

12.4 is the answer??

OpenStudy (anonymous):

damn...12m tires, that's one monster plow :)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok now im confused

OpenStudy (anonymous):

you have the circumference of the tire as 156/39=4 C=pi*d d=C/pi where d is diameter, C is circumerference, as khan said

OpenStudy (saifoo.khan):

It's 4. like i said. 4m is the circumfrence.

OpenStudy (saifoo.khan):



OpenStudy (anonymous):

oooh a cookie.....lol jk jk

OpenStudy (saifoo.khan):

Lmao. That's a Tyre. of 4m.

OpenStudy (albertrea):

ameryn, you are lovely, sehr schoon!

OpenStudy (anonymous):

oh ahhah thanks

OpenStudy (anonymous):

156 = 39 * ( pi * D) ->D = 156/ 39 pi = 1.27m